OpenStudy (anonymous):

solve using elimination 2x + 5y =19 6x-5y=-9

OpenStudy (anonymous):

I can eliminate the y's right now, right?

OpenStudy (smileyxl3):

6x-5y=-9 2x+5y=19 ---------- 8x=10 x=5/4 I think you can solve from there

OpenStudy (anonymous):

You would add the x's?

OpenStudy (smileyxl3):

No, now that y is eliminated, you would substitute x into the original equation